# arXiv:2603.01919v1[cs.CR] 02 Mar 2026

### 02 Mar 2026 v1

@@ Introduction
Frontier large language models (LLMs) have transformed numerous domains, becoming essential infrastructure for scientific research and daily applications [ECDGGGHKLLLMZM25, ZZLTWHMZZDDYCCJRLTLLNW23, BCEGHKLLLLNPRZ23]. By integrating advanced reasoning and domain knowledge, these models demonstrate enhanced effectiveness in complex tasks [WWSBIXCLZ22, O23, SATMWCSTCPPSGKSCMAWCMCGTLRBSKN23]. Given that many frontier LLMs are proprietary or require substantial computational resources, local deployment is often infeasible. In recent years, access to these models has been mediated through commercial APIs provided by companies like OpenAI and Google [promptrevolution2025unlocking, XWGZJ25, LBLTSYZNWKNYYZCMRAHZDLRRYWSOZYSKGCKHHCXSGHIZCWLMZK22]. However, official APIs often impose high pricing, payment barriers, and strict geographical restrictions [CZZ23]. For instance, the official OpenAI API cannot be accessed directly from several regions, such as China, Russia, and Iran [openai_supported_countries, openai_bloomberg]. As a result, a vast third-party provider market has emerged, offering compatible endpoints at a lower cost and without regional limitations. We refer to these services as shadow APIs (illustrated in Figure 1), platforms that claim to generate the same output as official LLMs via indirect access. By December 6, 2025, the most popular shadow APIs have accumulated 58,639 GitHub stars and have been utilized in 187 academic papers with 5,966 cumulative citations. Most of these papers have accepted by top venues, such as ACL, CVPR, and ICLR (see details in Section 3).
@@ 3. RQ3:
LLMs have rapidly become the backbone of modern natural language processing (NLP), machine learning (ML), and even computer vision (CV) research. A recent analysis of 16,193 papers shows that by 2024, over $60\%$ of NLP papers, around $20\%$ of ML papers, and nearly $10\%$ of CV papers are related to LLMs [XZLCLLWL25]. The number of LLM papers has exploded from a few hundred in 2019 to more than 7,000 in 2024. At the same time, access to frontier proprietary models is tightly controlled. Providers restrict API access to certain regions for legal and security reasons [google_gemini_regions]. For instance, Anthropic explicitly disallows sales in unsupported regions [anthropicUnsupportedRegions]. OpenAI likewise warns that accessing or reselling its API from unsupported countries may lead to account suspension [openaiUnsupportedCountries]. These restrictions collide with the geographic barriers of artificial intelligence (AI) research, given the reality that major AI conferences such as AAAI and CVPR attract large numbers of submissions and accepted papers from regions with restricted access to frontier APIs (e.g., China) [aaai26, latticeflowCVPR24Numbers].
In addition to geographic barriers, expensive bills put pressure on users. Frontier LLM APIs are typically priced for enterprise customers, which can be prohibitive for individual researchers or students. Analysis of LLM API economics suggests that many current usage patterns are unsustainable for small actors without cross-subsidies [tinymlLLMEconomics]. This drives the rise of a commercial shadow market offering discounted access to frontier LLMs without geographical restrictions [chinatalk2025grey, esentire2024genai].
Moreover, official terms of service prohibit any form of API key resale or redistribution [openai-services-agreement]. In addition, Chinese government regulations require AI services to operate in compliance with applicable laws and administrative requirements [cac_2023_measures]. As a result, these sellers operate in violation of both the service contract and applicable regulatory requirements. These trends suggest a growing disconnect between academic demand for frontier models and the constraints of official distribution channels, which drives the emergence of a substantial market of unofficial third-party services.
@@
We start from screening the accepted papers from ICLR 2024 (2,260 total, 1,108 with code) and papers from ACL 2024 (1,923 total, 1,005 with code) [iclr2024papers, acl2024papers]. To retrieve the associated code repositories, we systematically parsed the abstracts and footnotes of these papers to extract URLs matching standard repositories. From this combined pool of 2,113 code-available papers, we identify 92 projects that use LLM APIs via their associated GitHub repositories. Among these, 21 papers ($22.8\%$) use at least one shadow API endpoint, and we search GitHub for these URLs to find additional repositories that call the same endpoint URL, collecting the venue, institution, country, citation counts, and the GitHub stars of associated repositories, with all metrics updated as of December 6, 2025. We iterate this process by harvesting new shadow API endpoint URLs from these repositories, repeating until no new ones are discovered.
@@
Infrastructure. Among the 17 identified shadow API services, our analysis reveals that 11 are built upon open-source AI model aggregation and redistribution systems, primarily OneAPI [oneapi] and its derivative, NewAPI [newapi]. OneAPI is an open-source tool designed for self-hosted deployment. It unifies interfaces from various commercial LLM providers into a standard OpenAI-compatible format. The infrastructure system supports key features such as API key management, secondary redistribution, request routing, and automatic retries, thereby increasing the potential for exploitation, resale, and abuse more than the official API.
@@
Benchmarks and Methodology. We assess model utility across scientific and sensitive domains [T25, D25, T251]. For the science domain, we employ the AIME 2025 benchmark [aime2025], which tests competition-level mathematics, and the GPQA (Diamond) benchmark [RHSPPDMB23], targeting PhD-level scientific questions. For the sensitive domain, we focus on high-risk fields, i.e., medical and legal fields. In the medical field, following prior work [TSSCZXWZWSCG25], we use the MedQA (USMLE) dataset [JPOWFS21], covering diagnosis, treatment, and medical concepts. In the legal field, we select LegalBench [GNHRCNCPWRZTHSFSDPHWNNNCTHMLRHKHRGGWGZIL23], specifically the Scalr subset Rule-Application and Rule-Conclusion involving reasoning, consistent with [CTXWML25, HYGWKW25].
Implementation Details. To ensure fair comparison, we adopt EvalScope prompt templates [evalscope2025] for the AIME 2025 and GPQA benchmarks, and for the MedQA (USMLE) benchmark, we use the Hulu-Med multiple-choice template [JWSHZPZYFZHCWTLXWXFZLXSWL25], and the original task-specific prompts for LegalBench (Scalr) [GNHRCNCPWRZTHSFSDPHWNNNCTHMLRHKHRGGWGZIL23]. We obtain all results using the model configurations detailed in Table 6.
@@
Implementation Details. We employ four distinct jailbreak attacks to evaluate the safety: GCG [ZWCNKF23], Base64 [wei2023jailbroken], Combination [wei2023jailbroken], and FlipAttack [LHXFDH24]. For GCG, we use LLaMA3-8B to generate a universal suffix and then transfer it to other LLMs. For Base64 and Combination, we follow the settings for Base64 and combination_1 in [wei2023jailbroken]. For FlipAttack, we use its well-performed “flip char in sentence” mode.
Metrics. Following prior work such as PAIR [CRDHPW23], TAP [MZKNASK23], and JAIL-CON [jiang2025adjacent], we introduce a lightweight judge model to output the harmfulness score and verify whether the generated answer is valid and harmful. Specifically, based on [SLBTHPASEWT24], we build the evaluator with GPT-4o-mini and the rubric-based prompt template from StrongREJECT. A higher harmfulness score indicates a more harmful answer, i.e., lower safety.
@@ Conclusion
Detection Coverage and Ground Truth. In the absence of ground truth regarding the backend infrastructure of shadow APIs, our investigation relies on performance metrics, meta-information, and established auditing frameworks such as LLMmap and MET to detect model substitution and performance inconsistencies. While we can detect model identity mismatches with high confidence, we cannot definitively distinguish between certain fine-grained implementation details. Moreover, while we specify exact version snapshots for official models wherever possible, the official APIs of proprietary models may undergo undisclosed fine-tuning. To mitigate the potential impact of temporal drift and inherent stochasticity in official APIs, we conducted multiple independent runs for each experiment and reported the variance, ensuring that our comparative results remain robust against fluctuations in the baseline models.
@@ Appendix A Configurations Details
Table 6 details the specific configurations for all evaluated models across official and shadow APIs. To ensure reproducibility, we set the random seed at 42 and the temperature at 0 for all models compatible with these settings, in accordance with reasoning task guidelines [deepseek_params], while omitting these parameters for unsupported endpoints. For the GPT-5 model family and other reasoning models, we specifically monitor their support for reasoning parameters. For all reasoning models, we employ the medium reasoning effort in our subsequent analysis, according to the default configuration [openai_reasoning]. The table also highlights the discrepancies in which shadow APIs may enforce fixed behaviors that deviate from official baselines. The Ratio column denotes the multiplicative factor of the shadow API service price relative to the official provider for input/output tokens. Logprobs indicates whether the API can return log probability outputs. Pricing is normalized to USD per 1 million tokens for comparison and accessed on December 6, 2025.
@@
Shadow API A is selected as the first ranking in both categories, which is operated by an individual and does not employ open-source LLM API management or distribution systems. Shadow API E and H are selected for their stability, as they rank within the top for both citations and stars. Regarding shadow API E, the provider is also an individual and utilizes an open-source LLM API management and distribution system (NewAPI [newapi]) as its deployment template. Shadow API E aggregates over 54 upstream providers, some of which are explicitly labeled as having undisclosed origins. Some of the endpoints are categorized as reverse-engineered, web-based, capability-degraded, distilled, or unstable yet cost-effective variants. Given this diversity, we specifically select endpoints explicitly labeled as originating from the official. For the shadow API H, the provider is a corporate entity (Internet Content Provider filing) rather than an individual operator, which does not rely on an open-source LLM API management and distribution system. Notably, the shadow API B is deployed by a laboratory at a prestigious Chinese university, built on NewAPI, where part of the models listed in its internal large language Model Marketplace are explicitly marked with Unknown sources.

### 05 Mar 2026 v2

# Real Money, Fake Models: Deceptive Model Claims in Shadow APIs

@@ Introduction
Frontier large language models (LLMs) have transformed numerous domains, becoming essential infrastructure for scientific research and daily applications [ECDGGGHKLLLMZM25, ZZLTWHMZZDDYCCJRLTLLNW23, BCEGHKLLLLNPRZ23]. By integrating advanced reasoning and domain knowledge, these models demonstrate enhanced effectiveness in complex tasks [WWSBIXCLZ22, O232, SATMWCSTCPPSGKSCMAWCMCGTLRBSKN23]. Given that many frontier LLMs are proprietary or require substantial computational resources, local deployment is often infeasible. In recent years, access to these models has been mediated through commercial APIs provided by companies like OpenAI and Google [KS25, XWGZJ25, LBLTSYZNWKNYYZCMRAHZDLRRYWSOZYSKGCKHHCXSGHIZCWLMZK22]. However, official APIs often impose high pricing, payment barriers, and strict geographical restrictions [CZZ23]. For instance, the official OpenAI API cannot be accessed directly from several regions, such as China, Russia, and Iran [O252, B24]. As a result, a vast third-party provider market has emerged, offering compatible endpoints at a lower cost and without regional limitations. We refer to these services as shadow APIs (illustrated in Figure 1), platforms that claim to generate the same output as official LLMs via indirect access. By December 6, 2025, the most popular shadow APIs have accumulated 58,639 GitHub stars and have been utilized in 187 academic papers with 5,966 cumulative citations. Most of these papers have accepted by top venues, such as ACL, CVPR, and ICLR (see details in Section 3).
@@ 3. RQ3:
LLMs have rapidly become the backbone of modern natural language processing (NLP), machine learning (ML), as well as even computer vision (CV) research, and the base for building personal agents such as OpenClaw [JZSBZ26]. A recent analysis of 16,193 papers shows that by 2024, over $60\%$ of NLP papers, around $20\%$ of ML papers, and nearly $10\%$ of CV papers are related to LLMs [XZLCLLWL25]. The number of LLM papers has exploded from a few hundred in 2019 to more than 7,000 in 2024. At the same time, access to frontier proprietary models is tightly controlled. Providers restrict API access to certain regions for legal and security reasons [G25]. For instance, Anthropic explicitly disallows sales in unsupported regions [A25]. OpenAI likewise warns that accessing or reselling its API from unsupported countries may lead to account suspension [O24]. These restrictions collide with the geographic barriers of artificial intelligence (AI) research, given the reality that major AI conferences such as AAAI and CVPR attract large numbers of submissions and accepted papers from regions with restricted access to frontier APIs (e.g., China) [A252, L24].
In addition to geographic barriers, expensive bills put pressure on users. Frontier LLM APIs are typically priced for enterprise customers, which can be prohibitive for individual researchers or students. Analysis of LLM API economics suggests that many current usage patterns are unsustainable for small actors without cross-subsidies [A253]. This drives the rise of a commercial shadow market offering discounted access to frontier LLMs without geographical restrictions [Q252, K24].
Moreover, official terms of service prohibit any form of API key resale or redistribution [O251]. In addition, Chinese government regulations require AI services to operate in compliance with applicable laws and administrative requirements [CAC23]. As a result, these sellers operate in violation of both the service contract and applicable regulatory requirements. These trends suggest a growing disconnect between academic demand for frontier models and the constraints of official distribution channels, which drives the emergence of a substantial market of unofficial third-party services.
@@
We start from screening the accepted papers from ICLR 2024 (2,260 total, 1,108 with code) and papers from ACL 2024 (1,923 total, 1,005 with code) [I24, A24]. To retrieve the associated code repositories, we systematically parsed the abstracts and footnotes of these papers to extract URLs matching standard repositories. From this combined pool of 2,113 code-available papers, we identify 92 projects that use LLM APIs via their associated GitHub repositories. Among these, 21 papers ($22.8\%$) use at least one shadow API endpoint, and we search GitHub for these URLs to find additional repositories that call the same endpoint URL, collecting the venue, institution, country, citation counts, and the GitHub stars of associated repositories, with all metrics updated as of December 6, 2025. We iterate this process by harvesting new shadow API endpoint URLs from these repositories, repeating until no new ones are discovered.
@@
Infrastructure. Among the 17 identified shadow API services, our analysis reveals that 11 are built upon open-source AI model aggregation and redistribution systems, primarily OneAPI [Q25] and its derivative, NewAPI [Q251]. OneAPI is an open-source tool designed for self-hosted deployment. It unifies interfaces from various commercial LLM providers into a standard OpenAI-compatible format. The infrastructure system supports key features such as API key management, secondary redistribution, request routing, and automatic retries, thereby increasing the potential for exploitation, resale, and abuse more than the official API.
@@
Benchmarks and Methodology. We assess model utility across scientific and sensitive domains [T25, D25, T251]. For the science domain, we employ the AIME 2025 benchmark [A251], which tests competition-level mathematics, and the GPQA (Diamond) benchmark [RHSPPDMB23], targeting PhD-level scientific questions. For the sensitive domain, we focus on high-risk fields, i.e., medical and legal fields. In the medical field, following prior work [TSSCZXWZWSCG25], we use the MedQA (USMLE) dataset [JPOWFS21], covering diagnosis, treatment, and medical concepts. In the legal field, we select LegalBench [GNHRCNCPWRZTHSFSDPHWNNNCTHMLRHKHRGGWGZIL23], specifically the Scalr subset Rule-Application and Rule-Conclusion involving reasoning, consistent with [CTXWML25, HYGWKW25].
Implementation Details. To ensure fair comparison, we adopt EvalScope prompt templates [C25] for the AIME 2025 and GPQA benchmarks, and for the MedQA (USMLE) benchmark, we use the Hulu-Med multiple-choice template [JWSHZPZYFZHCWTLXWXFZLXSWL25], and the original task-specific prompts for LegalBench (Scalr) [GNHRCNCPWRZTHSFSDPHWNNNCTHMLRHKHRGGWGZIL23]. We obtain all results using the model configurations detailed in Table 6.
@@
Implementation Details. We employ four distinct jailbreak attacks to evaluate the safety: GCG [ZWCNKF23], Base64 [WHS23], Combination [WHS23], and FlipAttack [LHXFDH24]. For GCG, we use LLaMA3-8B to generate a universal suffix and then transfer it to other LLMs. For Base64 and Combination, we follow the settings for Base64 and combination_1 in [WHS23]. For FlipAttack, we use its well-performed “flip char in sentence” mode.
Metrics. Following prior work [CRDHPW23, MZKNASK23, JLBZ25, JHLZBZ26] such as PAIR, TAP, and JAIL-CON, we introduce a lightweight judge model to output the harmfulness score and verify whether the generated answer is valid and harmful. Specifically, based on [SLBTHPASEWT24], we build the evaluator with GPT-4o-mini and the rubric-based prompt template from StrongREJECT. A higher harmfulness score indicates a more harmful answer, i.e., lower safety.
@@ Conclusion
Detection Coverage and Ground Truth. In the absence of ground truth regarding the backend infrastructure of shadow APIs, our investigation relies on performance metrics, meta-information, and established auditing frameworks such as LLMmap and MET to detect model substitution and performance inconsistencies. While we can detect model identity mismatches with high confidence, we cannot definitively distinguish between certain fine-grained implementation details. Moreover, while we specify exact version snapshots for official models wherever possible, the official APIs of proprietary models may undergo undisclosed fine-tuning. To mitigate the potential impact of temporal drift and inherent stochasticity in official APIs, we conducted multiple independent runs for each experiment and reported the variance, ensuring that our comparative results remain robust against fluctuations in the baseline models. In addition, we recognize the potential to analyze the differences between official and shadow APIs from diverse perspectives, such as bias and hallucination [JLSLBZ24, Z251].
@@ Appendix A Configurations Details
Table 6 details the specific configurations for all evaluated models across official and shadow APIs. To ensure reproducibility, we set the random seed at 42 and the temperature at 0 for all models compatible with these settings, in accordance with reasoning task guidelines [D251], while omitting these parameters for unsupported endpoints. For the GPT-5 model family and other reasoning models, we specifically monitor their support for reasoning parameters. For all reasoning models, we employ the medium reasoning effort in our subsequent analysis, according to the default configuration [O25]. The table also highlights the discrepancies in which shadow APIs may enforce fixed behaviors that deviate from official baselines. The Ratio column denotes the multiplicative factor of the shadow API service price relative to the official provider for input/output tokens. Logprobs indicates whether the API can return log probability outputs. Pricing is normalized to USD per 1 million tokens for comparison and accessed on December 6, 2025.
@@
Shadow API A is selected as the first ranking in both categories, which is operated by an individual and does not employ open-source LLM API management or distribution systems. Shadow API E and H are selected for their stability, as they rank within the top for both citations and stars. Regarding shadow API E, the provider is also an individual and utilizes an open-source LLM API management and distribution system (NewAPI [Q251]) as its deployment template. Shadow API E aggregates over 54 upstream providers, some of which are explicitly labeled as having undisclosed origins. Some of the endpoints are categorized as reverse-engineered, web-based, capability-degraded, distilled, or unstable yet cost-effective variants. Given this diversity, we specifically select endpoints explicitly labeled as originating from the official. For the shadow API H, the provider is a corporate entity (Internet Content Provider filing) rather than an individual operator, which does not rely on an open-source LLM API management and distribution system. Notably, the shadow API B is deployed by a laboratory at a prestigious Chinese university, built on NewAPI, where part of the models listed in its internal large language Model Marketplace are explicitly marked with Unknown sources.